A baseball is thrown horizontally with an initial velocity of 20m/s from the height of 1.5m. How long will it take to reach the ground? How far will it go horizontally?


1
Expert's answer
2021-04-23T13:15:50-0400

Equations of motion of baseball are "y(t) = h - \frac{g t^2}{2}", "x(t) = v_0 t", where "h" is the initial height, and "v_0" is the initial horizontal velocity.

When the ball reaches the ground, "y(t) = 0", from where it takes "t_1 = \sqrt{\frac{2 h}{g}} \approx 0.55 s" to reach the ground.

Substituting the time "t_1" from the last equation into equation for "x(t)", obtain the horizontal distance covered: "s = x(t_1) = v_0 \sqrt{\frac{2 h}{g}} \approx 11.06 m"